# NoteLLM-2: Multimodal Large Representation Models for Recommendation

### v1

@@ 2.3 Performance of Multimodal Representation of Fine-tuned MLRMs
We conduct zero-shot experiments on several existing MLLMs in Appendix G, and find that zero-shot is insufficient to adapt MLLMs to representation tasks, performing inferiorly compared to a simple baseline, BM25 [40]. This is due to the misalignment between pre-training tasks and representation tasks [60, 19]. Hence, fine-tuning MLLMs for representation tasks is necessary.
We design three end-to-end MLRMs to examine the representation training method: MTomato-Base uses Tomato (our continually pre-trained LLM based on LLaMA 2 [47], which lacks vision perception ability) as the LLM, CLIP ViT-B [37] as the vision encoder and a randomly initialized Q-Former [25] as the connector for efficiency. MQwen-Base replaces Tomato with Qwen-Chat [2] in MTomato-Base. MQwen-bigG replaces CLIP ViT-B with ViT-bigG [17] in MQwen-Base. We set the visual embedding length of these models as $16$ for efficiency. For comparison, we choose two pre-trained MLLMs: BLIP-2 and Qwen-VL-Chat [3]. All vision encoders in these models are frozen for the larger batch size. In addition to MLRMs, we also compare other existing multimodal representation baselines. We provide details about baselines and MLRMs in Appendix E.
The results are shown in Table 1. Several interesting observations can be made. Firstly, MLRMs based on LLMs can outperform existing baselines significantly. Qwen-VL-Chat achieves a $10.78\%$ improvement in R@100 compared to the traditional method, METER Co-attn on all pairs. Secondly, the end-to-end training representation method can enhance the model’s multimodal representation ability. MQwen-bigG improves R@100 on all pairs by $6.31\%$ compared to Qwen-Chat. However, when the visual encoder is small, such as CLIP ViT-B, the enhancement in multimodal perception is not substantial. MTomato-Base only improves R@100 on all pairs by $1.31\%$ compared to Tomato, and MQwen-Base only improves R@100 on all pairs by $1.36\%$ compared to Qwen-Chat. Lastly, although MQwen-bigG is more efficient than Qwen-VL-Chat, which uses the same vision encoder and LLM (see Appendix E.3), there remains a performance gap.
@@
Table 2: Representation performance of fine-tuned MLRMs under different modal inputs (%).
@@
Multimodal 78.54 93.77 99.03 60.43 83.26 96.60 61.63 85.48 97.89
@@
We also evaluate all fine-tuned MLRMs under different modal inputs, as shown in Table 2. Firstly, we observe that MTomato-Base and MQwen-Base hardly represent the image content when the input consists solely of images, which shows ignoring visual information. Owing to its more robust vision encoder, MQwen-bigG exhibits superior vision representation capabilities. However, its performance still falls short when compared to BLIP-2 and Qwen-VL-Chat, both of which have undergone multimodal pre-training.
2.4 Exploring the Bias of Fine-tuned MLRMs
@@
To further mine the bias in fine-tuned MLRMs, we explore the attention patterns of two modality tokens in different LLMs’ layers. Following [48], we adopt saliency scores to analyze the interaction between the compressed word and two modality tokens in the attention matrices. We randomly select $10,000$ notes in batches from the training dataset to count the saliency scores, and compute the saliency scores as follows:
$I_{l}=\sum_{h}\left|A^{\top}_{h,l}\frac{\partial\mathcal{L}}{\partial A_{h,l}}\right|,$ (3)
@@
where $A_{h,l}$ means the attention matrix value of the $h$-th attention head in the $l$-th layer and $\mathcal{L}$ is the contrastive loss. We sum all attention heads to obtain the saliency matrix $I_{l}$ of $l$-th layer. $I_{l}(i,j)$ means the significance of the information flow from the $j$-th word to the $i$-th word. To analyze different modality importances for the compressed word, we separate $I_{l}$ into three parts as flowing: $S_{v}$, the mean significance of information flow from the vision embeddings $E_{v}$ to the final compressed position $c$. $S_{t}$, the mean significance of information flow from the textual embeddings $E_{t}$ to the final compressed position $c$. $S_{o}$, the mean significance of information flow amongst all words, excluding influences represented by $S_{v}$ and $S_{t}$. We provide a more clear formal definition in Appendix F.1.
@@
The results of three parts from the saliency matrix are shown in Figure 3. MLRMs trained end-to-end without pre-training exhibit a low $S_{v}$ and a high $S_{o}$ in shallow layers, while pre-trained MLLMs, BLIP-2 and Qwen-VL-Chat (see Appendix F.2), have more balanced $S_{v}$ and $S_{o}$ in shallow layers. This indicates that end-to-end trained MLRMs primarily utilize the information gain from multimodal data to enhance the information flow in the context. Images in these models primarily assist the information flow in identifying crucial information [44, 43], rather than being directly aggregated by representations. Besides, text information flow significantly surpasses visual information flow in end-to-end trained MLRMs.
((a)) BLIP-2
((b)) MTomato-Base
((c)) MQwen-Base
((d)) MQwen-bigG
Figure 3: Relative sizes of $S_{v}$, $S_{t}$ and $S_{o}$ in different layers of different MLRMs.
3 Methodology
@@
From our investigation, MLRMs trained end-to-end without pre-training are suboptimal, primarily due to the disregard of visual information after processing by LLMs. Therefore, we aim to design mechanisms that focus more on visual signals. We propose a novel training framework, NoteLLM-2, which contains two methods from different perspectives. The first one is from the prompt viewpoint, called mICL. This method alters the prompt to adjust the attention patterns towards visual information. The second one is from the model architecture perspective, which combines the late fusion with the visual prompt. This enhances the impact of the visual information flow on the final representation by delaying the fusion of visual information. The improved overall framework is shown in Figure 2 (b).
@@
Specifically, given note $n_{i}$, the mICL mechanism does not attempt compress the multimodal information into one compressed word. Instead, we separate the multimodal note into two single-modality notes. We then adopt a similar ICL way [12, 48] to aggregate the multimodal information. We reformulate our note compression prompt as follows: {mdframed} mICL Note Compression Prompt.
@@
Note content: {’image’: <IMG>}, Compress this note into one word:"<IMG_EMB>". Note content: {’title’: $t_{i}$ , ’topic’: $tp_{i}$ , ’content’: $ct_{i}$}. Compress this note into one word:" In this template, <IMG_EMB> is a special token.
@@
After processing the multimodal embeddings with LLMs, we choose relevant hidden states to represent notes. Following [60, 19], we take the previous hidden state of the token <IMG_EMB> as the visual note representation, denoted as $\boldsymbol{n}_{i}^{v}$. Due to the causal attention, $\boldsymbol{n}_{i}^{v}$ only contains the note image information. We also take the last hidden state as the multimodal note representation, denoted as $\boldsymbol{n}_{i}^{m}$, which can compress visual and textual information. This method separates the original single embedding into two modality note embeddings. The multimodal note representation $\boldsymbol{n}_{i}^{m}$ aggregates useful visual information in a similar manner to ICL.
@@
The late fusion mechanism adopts original visual embedding to enhance the note embeddings. This prevents a textual bias due to the space of LLMs [62] and incorporates more original visual information. Specifically, after extracting image features by the vision encoder $V_{\theta}$, we take the visual feature, which contains the entire image information, such as the [CLS] token in CLIP ViT-B. Then, we use a linear layer to transform these features into the dimension of LLMs, denoted as $\boldsymbol{v}\in\mathbb{R}^{h_{t}}$. We adopt the same gate mechanism to fuse the original visual information into the two note embeddings.
$\boldsymbol{z}$ $=sigmoid(\boldsymbol{W}[\boldsymbol{v},\boldsymbol{n}_{i}^{v}]+\boldsymbol{b}),$ (4)
@@
$\boldsymbol{z}$ $=sigmoid(\boldsymbol{W}[\boldsymbol{v},\boldsymbol{n}_{i}^{m}]+\boldsymbol{b}),$ (5)
$\hat{\boldsymbol{n}_{i}^{m}}$ $=\boldsymbol{z}\odot\boldsymbol{v}+(\boldsymbol{1}-\boldsymbol{z})\odot\boldsymbol{n}_{i}^{m},$
where $\hat{\boldsymbol{n}_{i}^{v}}$ and $\hat{\boldsymbol{n}_{i}^{m}}$ means the fused note embeddings. [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] means the concatenate operation, and $\boldsymbol{W}$ and $\boldsymbol{b}$ are learnable parameters. $\odot$ is the element-wise product.
Next, we adopt the two fused embeddings to conduct contrastive learning as follows:
$\mathcal{L}^{v}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{i}^{v,+}})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{j}^{v}})\cdot e^{\tau}}},$ (6)
$\mathcal{L}^{m}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\hat{\boldsymbol{n}_{i}^{m}},\hat{\boldsymbol{n}_{i}^{m,+}})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\hat{\boldsymbol{n}_{i}^{m}},\hat{\boldsymbol{n}_{j}^{m}})\cdot e^{\tau}}},$ (7)
where $\mathcal{L}^{v}(\pi,\mu)$ is the loss from visual note embeddings and $\mathcal{L}^{m}(\pi,\mu)$ is the loss from the multimodal note embeddings. The final loss is computed as follows:
$\mathcal{L}^{f}(\pi,\mu)=\frac{\mathcal{L}^{v}(\pi,\mu)+\alpha\mathcal{L}^{m}(\pi,\mu)}{1+\alpha},$ (8)
@@
4 Experiments
Table 3: Performance of NoteLLM-2 based on three different MLRMs (%).
  Method All Pair Short Query Pair Short Target Pair

### v2

@@ 2.4. Performance of Multimodal Representation of Fine-tuned MLRMs
We conduct zero-shot experiments on several existing MLLMs in Appendix D, and find that zero-shot is insufficient to adapt MLLMs to representation tasks, performing inferiorly compared to a simple baseline, BM25 (Robertson et al., 2009). This is due to the misalignment between pre-training tasks and representation tasks (Zhang et al., 2024a; Jiang et al., 2023a). Hence, fine-tuning MLLMs for representation tasks is necessary.
@@
Table 3. Performance of MLRMs fine-tuned with multimodal inputs under different modal inputs (%).
@@
Multimodal 78.53 93.76 99.03 60.41 83.27 96.60 61.60 85.48 97.89
We design three end-to-end MLRMs to examine the representation fine-tuning method: MTomato-Base uses Tomato (our continually pre-trained LLM based on LLaMA 2 (Touvron et al., 2023), which lacks vision perception ability. See Appendix B for more details on pre-training.) as the LLM, CLIP ViT-B (Radford et al., 2021) as the vision encoder and a randomly initialized Q-Former (Li et al., 2023) as the connector for efficiency. MQwen-Base replaces Tomato with Qwen-Chat (Bai et al., 2023a) in MTomato-Base. MQwen-bigG replaces CLIP ViT-B with ViT-bigG (Ilharco et al., 2021) in MQwen-Base. We set the visual embedding length of these models as $16$ for efficiency. For comparison, we choose two pre-trained MLLMs: BLIP-2 and Qwen-VL-Chat (Bai et al., 2023b). All vision encoders in these models are frozen for the larger batch size. We provide details about end-to-end MLRMs in Appendix B.
To evaluate the effectiveness of MLRMs, we compare them with several baselines. (1) CLIP ViT-B (Radford et al., 2021) is trained using only image input. (2) BM25 (Robertson et al., 2009) is a basic baseline, which counts the relevance between the query and documents based on every word in the query. (3) RoBERTa-wwm-ext (Cui et al., 2021) is a classic traditional Chinese BERT-like (Kenton and Toutanova, 2019) text encoder with pure text input. (4) Tomato is our continually pre-trained LLM that is based on LLaMA 2 (Touvron et al., 2023). (5) NoteLLM (Zhang et al., 2024a) unifies the embedding generation and topic generation into one task. (6) METER Merge-attn and METER Co-attn (Dou et al., 2022) leverage best training strategy to train vision-and-language transformers. METER Merge-attn adopts the merged attention as the fusion module, while METER Co-attn uses co-attention as the fusion module. All baselines are trained on our training dataset.
The results are shown in Table 2. Several observations can be made. Firstly, MLRMs can significantly outperform existing non-LLM-based baselines. Qwen-VL-Chat achieves a $10.78\%$ improvement in R@100 compared to the traditional method, METER Co-attn on all pairs. Secondly, the end-to-end fine-tuned representation method can enhance the model’s multimodal representation ability. MQwen-bigG improves R@100 on all pairs by $6.31\%$ compared to Qwen-Chat. However, when the visual encoder is small, such as CLIP ViT-B, the enhancement in multimodal perception is not substantial. MTomato-Base only improves R@100 on all pairs by $1.31\%$ compared to Tomato, and MQwen-Base only improves R@100 on all pairs by $1.36\%$ compared to Qwen-Chat. Lastly, although MQwen-bigG is more efficient than Qwen-VL-Chat, which uses the same vision encoder and LLM (see Section 4.4), there remains a performance gap.
@@
We also evaluate all MLRMs fine-tuned with multimodal inputs under different modal inputs, as shown in Table 3. Firstly, we observe that MTomato-Base and MQwen-Base hardly represent the image content when the input consists solely of images, which ignores visual information. Owing to its more robust vision encoder, MQwen-bigG exhibits superior vision representation capabilities. However, its performance still falls short when compared to BLIP-2 and Qwen-VL-Chat, both of which have undergone multimodal pre-training.
2.5. Exploring the Information Flow of Fine-tuned MLRMs
@@
To further explore the reasons why end-to-end fine-tuned MLRMs overlook visual information, we examine the attention patterns of two modality tokens in different LLMs’ layers from the perspective of information flow. We adopt saliency scores (Wang et al., 2023b) to analyze the interaction between the compressed word and two modality tokens in the attention matrices. We randomly select $10,000$ notes in batches from the training dataset to count the saliency scores, and compute the saliency scores as follows:
(3) $I_{l}=\sum_{h}\left|A^{\top}_{h,l}\frac{\partial\mathcal{L}}{\partial A_{h,l}}\right|,$
@@
where $A_{h,l}$ means the attention matrix value from the $h$-th attention head within $l$-th layer and $\mathcal{L}$ is the contrastive loss. The saliency matrix $I_{l}$ for layer $l$ is computed by summing across all attention heads. $I_{l}(i,j)$ means the significance of the information flow from the $j$-th word to the $i$-th word. To analyze the relative importance of different modalities for the compressed word, we separate $I_{l}$ into three parts as follows: $S_{v}$, the mean significance of information flow from the vision embeddings $E_{v}$ to the final compressed position $c$:
(4) $S_{v}$ $=\frac{\sum_{(i,j)\in P_{v}}I_{l}(i,j)}{\left|P_{v}\right|},$
$P_{v}$ =$=\{(c,j):j\in E_{v}\}.$
(5) $S_{t}$ $=\frac{\sum_{(i,j)\in P_{t}}I_{l}(i,j)}{\left|P_{t}\right|},$
$P_{t}$ =$=\{(c,j):j\in E_{t},j<c\}.$
(6) $S_{o}$ $=\frac{\sum_{(i,j)\in P_{o}}I_{l}(i,j)}{\left|P_{o}\right|},$
$P_{o}$ =$=\{(i,j):j<i\}-P_{v}-P_{t}.$
@@
The results of three parts from the saliency matrix are shown in Figure LABEL:fig:saliency_score. We observe that pre-trained MLLMs, BLIP-2 and Qwen-VL-Chat (see Appendix C), have balanced $S_{v}$ and $S_{o}$ in shallow layers. However, MLRMs fine-tuned end-to-end without pre-training exhibit a low $S_{v}$ and a high $S_{o}$ in shallow layers. The low $S_{v}$ indicates limited direct information flow from visual to final compressed embeddings, while the high $S_{o}$ indicates significant information flow between visual and textual embeddings. This shows that the information flow gain brought by introducing images is more likely to flow towards important multimodal embeddings, rather than being directly aggregated into final compressed embeddings. As the number of layers increases, we observe that $S_{v}$ increases while $S_{o}$ decreases dramatically. This is because, after multimodal embeddings capture context information in shallow layers (as shown by $S_{o}$), models focus on directly aggregating relevant information into final compressed embeddings for representation tasks.
Unlike pre-trained MLLMs, the final compressed embeddings of end-to-end MLRMs have little direct visual information in shallow layers. This causes the final compressed embedding initially accumulates visual information mainly through the aggregation of context information contained in textual embeddings. In deep layers, final compressed embeddings increasingly incorporate image information related to the previously aggregated textual information through the attention mechanism. Therefore, images in end-to-end MLRMs primarily assist the information flow in identifying crucial information (Sun et al., 2021, 2020), rather than being directly aggregated by representations. Besides, text information flow significantly surpasses visual information flow in end-to-end fine-tuned MLRMs.
Table 4. Performance of different methods based on three MLRMs (%). “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
  Base Model Method All Pair Short Query Pair Short Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
  MTomato-Base basic method 71.94 88.22 96.13 44.77 63.77 83.31 45.83 66.08 87.02
mICL 74.15 90.94 98.26 51.56 75.87 93.84 52.74 78.13 95.82
late fusion 74.51 91.37 98.45 54.19 78.02 94.52 54.00 79.71 95.80
NoteLLM-2 74.62* 91.50* 98.47 54.87* 78.82* 94.69* 54.51* 79.75 95.46
only late fusion 74.23 91.25 98.43 53.20 77.75 94.65 53.65 79.76 95.59
Omni-Retrieval 71.73 88.12 96.32 44.98 64.26 84.57 45.18 66.14 88.48
  MQwen-Base basic method 74.02 89.65 97.22 48.15 68.35 88.22 49.82 70.91 91.43
mICL 75.81 91.57 98.55 53.32 76.40 94.38 54.55 79.32 96.05
late fusion 75.92 91.50 98.28 52.74 76.03 93.65 54.06 78.89 95.53
NoteLLM-2 76.49* 92.18* 98.57 55.73* 78.91* 94.94* 56.12* 80.65* 96.24*
only late fusion 75.84 91.48 98.35 53.33 76.22 93.57 54.27 78.38 95.29
Omni-Retrieval 74.71 91.05 98.11 51.61 74.34 92.90 51.91 77.07 94.71
  MQwen-bigG basic method 77.63 92.89 98.91 57.45 80.92 95.73 57.95 83.50 96.99
mICL 77.96* 93.18 98.92 58.48* 82.09* 96.65* 58.76 84.01 97.20*
late fusion 77.43 92.93 98.78 57.47 81.78 95.89 58.48 83.99 97.16
NoteLLM-2 77.56 93.25 98.82 57.94 81.98 96.29 58.76 83.87 97.05
only late fusion 76.23 92.30 98.67 55.21 79.22 95.58 55.16 81.13 96.70
Omni-Retrieval 77.38 93.11 98.91 58.26 82.07 96.61 58.28 83.95 97.13
3. Methodology
@@
From our investigation, MLRMs trained end-to-end without pre-training are suboptimal, primarily due to the disregard of visual information after processing by LLMs. Therefore, we aim to design mechanisms that focus more on visual signals. We propose a novel training framework, NoteLLM-2, which contains two methods from different perspectives. The first one is from the prompt viewpoint and is called mICL. This method alters the prompt to adjust the attention patterns towards visual information. The second one is from the model architecture perspective, which combines the late fusion with the visual prompt. This enhances the impact of the visual information flow on the final representation by delaying the fusion of visual information. The improved overall framework is shown in Figure 2 (b).
@@
Specifically, given note $n_{i}$, the mICL does not compress the multimodal information into one compressed word. Instead, we separate the multimodal note into two single-modality notes. We then adopt a similar ICL way (Dong et al., 2022; Wang et al., 2023b) to aggregate the multimodal information. We reformulate our note compression prompt as follows:
{mdframed}
mICL Note Compression Prompt.
@@
Note content: {‘image’: ¡IMG¿}, Compress this note into one word:“¡IMG_EMB¿”. Note content: {‘title’: $t_{i}$ , ‘topic’: $tp_{i}$ , ‘content’: $ct_{i}$}. Compress this note into one word:“ In this template, ¡IMG_EMB¿ is a special token.
@@
After processing the multimodal embeddings with LLMs, we choose relevant hidden states to represent notes. Following (Zhang et al., 2024a; Jiang et al., 2023a), we take the previous hidden state of the token ¡IMG_EMB¿ as the visual note representation, denoted as $\boldsymbol{n}_{i}^{v}$. Due to the causal attention, $\boldsymbol{n}_{i}^{v}$ only contains the note image information. We also take the last hidden state as the multimodal note representation, denoted as $\boldsymbol{n}_{i}^{m}$, which contains visual and textual information. This method separates the original single embedding into two modality note embeddings. The multimodal note representation $\boldsymbol{n}_{i}^{m}$ aggregates useful visual information in a similar manner to ICL.
@@
The late fusion adopts original visual embedding to enhance the note embeddings. This prevents a textual bias due to the space of LLMs (Zhang et al., 2024b) and incorporates more original visual information. Specifically, after extracting image features by the vision encoder $V_{\theta}$, we take the visual feature, which contains the entire image information, such as the [CLS] token in CLIP ViT-B. Then, we use a linear layer to transform these features into the dimension of LLMs, denoted as $\boldsymbol{v}\in\mathbb{R}^{h_{t}}$. We adopt the gate mechanism (Chung et al., 2014) to fuse the original visual information into the note embeddings.
(7) $\boldsymbol{z}$ $=sigmoid(\boldsymbol{W}[\boldsymbol{v},\boldsymbol{n}_{i}^{v}]+\boldsymbol{b}),$
@@
where $\hat{\boldsymbol{n}_{i}^{v}}$ means the fused visual note embeddings. [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] means the concatenate operation, and $\boldsymbol{W}$ and $\boldsymbol{b}$ are learnable parameters. $\odot$ is the element-wise product. Similarly, we can get the fused multimodal note embedding $\hat{\boldsymbol{n}_{i}^{m}}$ with the same gate operation.
Next, we adopt the fused note embeddings to conduct contrastive learning as follows:
(8) $\mathcal{L}^{v}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{i}^{v,+}})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{j}^{v}})\cdot e^{\tau}}},$
where $\mathcal{L}^{v}(\pi,\mu)$ is the loss from visual note embeddings. Similarly, we can get $\mathcal{L}^{m}(\pi,\mu)$, which is the loss from the multimodal note embeddings. The final loss is computed as follows:
(9) $\mathcal{L}^{f}(\pi,\mu)=\frac{\mathcal{L}^{v}(\pi,\mu)+\alpha\mathcal{L}^{m}(\pi,\mu)}{1+\alpha},$
@@
4. Experiments
4.1. Performance Evaluation
We conduct experiments on following baselines based on three MLRMs to show NoteLLM-2’s effectiveness. basic method is the naive representation method of MLRMs. mICL adds the mICL method based on the basic method. late fusion integrates the late fusion mechanism into the basic method. NoteLLM-2 contains both mICL and late fusion simultaneously. only late fusion only uses late fusion to integrate image and text information, without inputting image embeddings into LLMs. Omni-Retrieval (Yu et al., 2022a) adopts cross-modal contrastive learning across various embeddings, including image-only, text-only, and multimodal, to achieve alignment.
The results are presented in Table 4. Firstly, NoteLLM-2 can significantly enhance the overall performance of MTomato-Base and MQwen-Base, which have relatively small vision encoders. Additionally, NoteLLM-2 primarily enhances MQwen-bigG in terms of short pairs. Secondly, mICL enhances the performance of all models, while late fusion is more effective in models with relatively small vision encoders. When the vision encoders are powerful, MQwen-bigG+late fusion improves performance on short pairs by focusing more on images. However, this overemphasis on images can detrimentally affect text comprehension, reducing overall performance and weakening the mICL method that utilizes LLM prompts. We will design a more balanced late fusion mechanism in future work. Concurrently, only late fusion is a straightforward and efficient fusion method. However, its simplicity could potentially be a drawback when the vision encoder is more powerful, as it may not fully interact with LLMs, leading to a loss in performance. Besides, the enhancement of Omni-Retrieval for visual information is still limited compared to NoteLLM-2. This is because Omni-Retrieval ignores the distinct mechanisms of LLMs, and other cross-modal losses may interfere with multimodal embeddings. Lastly, MQwen-bigG+NoteLLM-2 underperforms compared to Qwen-VL-Chat, especially in shorter pairs. This is due to the visual token length difference between MQwen-bigG ($16$) and Qwen-VL-Chat ($256$).
4.2. Saliency Scores of Enhanced MLRMs
To further understand the impact of NoteLLM-2 on MLRMs, we illustrate the difference in saliency scores of enhanced MLRMs. The saliency scores for the original fine-tuned method are $S_{v}$, $S_{t}$, and $S_{o}$. The saliency scores for the enhanced fine-tuned method are $\hat{S_{v}}$, $\hat{S_{t}}$ and $\hat{S_{o}}$. We consider the visual note compressed word as a part of the vision embeddings $E_{v}$. Our results are shown in Figure LABEL:fig:ssdif. The representations of all enhanced MLRMs amplify $S_{v}$, the direct focus on images, while reducing the emphasis on $S_{o}$ in the shallow layers. At the same time, $S_{t}$ remains virtually unchanged. More balanced $S_{v}$ and $S_{o}$ mean that the information flow gain brought by introducing images not only highlights the relevant multimodal embeddings but is also aggregated by the final compressed embedding directly. This is helpful for final compressed embeddings to capture crucial patterns in images that are totally different from text. This result is attributed to mICL’s use of the same compressed prompt to compress both modalities. By identifying similar compression patterns from images, mICL can bolster the concentration of multimodal representations on images, similar to ICL (Wang et al., 2023b).
4.3. Hyper-Parameter Analysis
In this section, we use MTomato-Base+NoteLLM-2 to conduct hyper-parameter analysis experiments.
Impact of length of visual tokens. We explore the impact of visual token length. The results are shown in Table 5. Reducing the length from $16$ to $8$ leads to poorer performance, while inference is more efficient. There exists a tradeoff between performance and inference speed concerning the length of visual tokens. Moreover, MTomato-Base+NoteLLM-2 with 8 visual tokens outperforms MTomato-Base with 16 visual tokens in Table 2. This shows that our method enhances visual attention by modified attention patterns and fusion, rather than merely expanding image token space.
Table 5. Impact of the number of visual tokens $L_{c}$ (%). Inf. Sp. means the inference speed.
  $L_{c}$ Inf. Sp. All Pair Short Query Pair Short Target Pair